OpenStudy (anonymous):

Simplify the expression 3 over quantity of 5x to the negative second power. options: A.75x2 B.30x2 C.3 x squared over 5 D.25 x squared over 3

OpenStudy (anonymous):

i think the answer is D am i right

OpenStudy (anonymous):

3/(5x^-2) = 3(5x)^2 = 3(25x^2) = 75x^2 So your answer will be A

OpenStudy (zzr0ck3r):

\((\frac{3}{5x})^{-2}=(\frac{5x}{3})^2=\frac{(5x)^2}{3^2}=\frac{5^2x^2}{9}=\frac{25x^2}{9}\)

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

\((\frac{3}{5x})^{-2}\) or \(\frac{3}{(5x)^{-2}}\)?

OpenStudy (zzr0ck3r):

then yes A

OpenStudy (anonymous):

3/(5x)−2

OpenStudy (anonymous):

sorry this one

OpenStudy (zzr0ck3r):

\(\frac{3}{(5x)^{-2}}=3*(5x)^2=3*5^2*x^2=75x^2\)

OpenStudy (anonymous):

ok thanks